OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 3.0 centimeters/minute while the area of the triangle is increasing at a rate of 3.5 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 9.5 centimeters and the area is 94.0 square centimeters?

OpenStudy (gorv):

rate of change or area= rate of change of base* rate of change of altitude

OpenStudy (anonymous):

So how would I get that? (94-3.5*9.5-3.0)?

OpenStudy (anonymous):

57.75 that's it

OpenStudy (phi):

This is a "related rates" problem in calculus. Start with the formula for the area of a triangle \[ A = \frac{1}{2} b\ h \] Using the numbers given, A= 94 \( cm^2\) and h= 9.5 cm we can solve for the base \[ b = 2A/h= 19.79 \ cm \] Next, find the derivative with respect to time (variable "t" ). We assume all the variables A, b and h are changing. I will use A' to show dA/dt, b' for db/dt, etc Using the product rule we find \[ \frac{d}{dt} \left(A = \frac{1}{2} b\ h \right) \\ A' =\frac{1}{2} \left( b \ h' + h \ b'\right) \\ 2A' = b \ h' + h \ b' \] Fill in the variables that we know. altitude of a triangle is increasing at a rate of 3.0 centimeters/minute i.e. h' = + 3 cm/min area of the triangle is increasing at a rate of 3.5 square centimeters/minute. A' = + 3.5 cm^2 / min the altitude is 9.5 centimeters h = 9.5 cm also, from the first step, b= 19.79 cm we have \[ 2 \cdot 3.5 = 19.79 \cdot 3+ 9.5 \cdot b' \\ 7 = 59.37 + 9.5 b' \\ -52.37 / 9.5 = b' \\ b'= - 5.5\ cm/min \] the base is getting shorter at a rate of 5.5 cm/min